a. A student dissolved 20.00 g of Na2SO4 in enough water to make 250 mL of stock solution. How many mL of the stock solution would the student need to use in order to make a 150 mL of 0.100 M solution?

b. In an acid-base titration, 82.56 mL of 0.175 M NaOH titrant reacts with 100.0 mL of an unknown concentration of HCl solution. What is the density of the HCl solution?  
Answer #1

ANSWER:-

A)- GIVEN WEIGHT=20.0gm of Na2SO4 and volume is 250 ml molar mass of Na2SO4 is 142g

so first you have to calculate the molarity of Na2SO4 solution from the formula

molarity M = w * 1000/m * v

M= 20g * 1000(ml)/142g * 250 ml

M1= 0.56 of (Na2SO4) then use formula

M1V1 = M2V2

0.56 * V1 = 150 * 0.100

V1= 26.78 mL

SO, 26.78 ml of stock solution will use.

B)- FOR CALCULATE THE DENSITY WE SHOULD FIRST CALCULATE THE MOLARITY OF HCL SOLUTION USING FORMULA,

M1V1 = M2V2

82.56 ml (NaOH) * 0.175M = 100 ml(HCL) * M2

MOLARITY OF HCL =0.144 M

then we use molarity formula for calculating density

molarity = w * 1000/m * v

and we know that density = m/v

so from above formula

molarity = d *1000/m

and density = molarity * m/1000

where m(molar mass) of HCL = 36.45

so density d= 0.144M * 36.5(g)/ 1000(ml)

(density of HCL) d= 0.005256 g/ml
Answer #2

ANSWER:-

A)- GIVEN WEIGHT=20.0gm of Na2SO4 and volume is 250 ml molar mass of Na2SO4 is 142g

so first you have to calculate the molarity of Na2SO4 solution from the formula

molarity M = w * 1000/m * v

M= 20g * 1000(ml)/142g * 250 ml

M1= 0.56 of (Na2SO4) then use formula

M1V1 = M2V2

0.56 * V1 = 150 * 0.100

V1= 26.78 mL

SO, 26.78 ml of stock solution will use.

B)- FOR CALCULATE THE DENSITY WE SHOULD FIRST CALCULATE THE MOLARITY OF HCL SOLUTION USING FORMULA,

M1V1 = M2V2

82.56 ml (NaOH) * 0.175M = 100 ml(HCL) * M2

MOLARITY OF HCL =0.144 M

then we use molarity formula for calculating density

molarity = w * 1000/m * v

and we know that density = m/v

so from above formula

molarity = d *1000/m

and density = molarity * m/1000

where m(molar mass) of HCL = 36.45

so density d= 0.144M * 36.5(g)/ 1000(ml)

(density of HCL) d= 0.005256 g/ml